A company is saving for some new equipment. They plan to invest $18,034 per year starting the end of this year, and increasing the investment by 4.3% per year. If the interest rate is 4.3%, how much can they spend in 7 years?
Answer #1

i=g=4.3% = 0.043

n=7 years

PV of the geometric series when i=g is given by P=C*n/(1+i)

= 18034*7/(1+0.043)

= 126238/1.043

= 121033.6